# SIGMMA: Hierarchical Graph-Based Multi-Scale Multi-modal Contrastive Alignment of Histopathology Image and Spatial Transcriptome

### v1

@@ 6 Experiments
We conduct extensive benchmarking on the HEST-1k dataset [jaume2024hest], the largest publicly available dataset of paired HE and ST data. Four subsets of the dataset provide paired HE–Xenium ST data, covering four cancer types: Invasive Ductal Carcinoma (IDC), Pancreatic Adenocarcinoma (PAAD), Skin Cutaneous Melanoma (SKCM), and Lung Adenocarcinoma (LUAD). In addition to the public datasets, we include an in-house skin dataset. Following the tiling scheme in multiple histopathology image foundation models [chen2024towards, vorontsov2024foundation, hoptimus0], we tessellate each WSI into $256\times 256$ pixel-sized tiles at 20x magnification level, which corresponds to $0.5\mu\text{m/pixel}$ resolution. For more details on data preprocessing/data splits, see Supplementary Material.
@@
We comprehensively compare sigmma against three categories of baselines: (1) Uni-modal HE image encoder, UNI [chen2024towards]; (2) Multi-modal vision–language (VL) models pre-trained on natural image–caption pairs, CLIP [radford2021learning], or medical text, PLIP [huang2023visual]; (3) HE–ST contrastive alignment models, including OmiCLIP [chen2025visual], which uses a text encoder by representing ST as a 1D sequence of gene names, and BLEEP [xie2023spatially], a simple MLP operating on batch corrected ST principal components. To ensure fair comparison, all baselines are fine-tuned on the datasets used in this study. For each method, we adopt the hyperparameters reported in the original paper; when unavailable, we determine them through grid search. For more details on the experiment setting, see Supplementary Material.
@@
Here, our focus is to evaluate the quality of the learned image representation for gene expression prediction. Therefore, to avoid introducing biases from different methods’ gene expression decoders, we use the image embedding output by each method and apply a ridge regression for gene expression prediction for each method. Multi-modal alignment consistently enriches HE image embeddings by incorporating ST information (Tab. 3). When applying sigmma on top of ResNet50 [he2016deep], H-Optimus-0 [hoptimus0], or UNI [chen2024towards], Sigmma improves representations across backbones, achieving up to 67% lower MSE and 56% higher PCC with UNI. Given its strong gains, we used UNI as the HE encoder backbone for all subsequent experiments.
Table 3: Ablation of vision backbone in HEST1k-LUAD dataset.
Task 1. Gene expression prediction.
Model MSE (↓) PCC (↑)
[…]
@@ 6.4 Ablation study
As shown in Tab. 4, we analyze the impact of the core components of sigmma: (1) cell graph, (2) multi-scale loss, and (3) graph sparsification via stochastic edge addition. For the ablation of the cell graph, we replaced the spatial graph representation with a 1D sequence of genes. For the ablation of multi-scale loss, we removed the micro- and meso-scale objectives and trained only with the macro-scale (single-scale) alignment loss. For the ablation of graph sparsification module, instead of selectively sampling edges through stochastic addition, we connected all nodes within neighboring patches, resulting in a fully connected local graph. Tab. 4 shows that each component contributes to performance gains: adding the multi-scale loss and graph sparsification progressively improves prediction accuracy, with all components combined achieving the lowest MSE and highest PCC. The ablation study for Task 2 is provided in the Supplementary Material. Overall, the ablation study shows that each component contributes to performance, with multi-scale loss and graph sparsification having the largest impact.
@@ 6.5 Biological application
We next evaluated whether the multiscale cell-level embeddings learned by sigmma capture meaningful biological structure in a publicly available section from PAAD tissue. For more details on the section, see Supplementary Material, which includes a full description of the dataset composition and sources. We performed clustering of sigmma embeddings at the micro, meso, and macro-scale and projected these clusters back onto a PAAD tissue section, comparing them with reference cell labels (Fig. 4a). Across all three scales, the inferred microenvironments formed coherent and spatially contiguous domains. Most notably at the micro- and meso-scales, sigmma cleanly delineated tumor nests from surrounding non-tumor tissue, despite the absence of any supervision from cell-type labels. To characterize the cellular context represented by these embeddings, we quantified the cell-type composition of each microenvironment cluster, focusing on the micro-scale (Fig. 4b). We identified six resolvable microenvironments. Two microenvironments were composed predominantly of tumor cells. A third represented an inflammatory infiltrate enriched for multiple immune cell types, including T cells and B cells, that were spatially excluded from the tumor region. Three additional microenvironments corresponding to known pancreatic structures, including perivascular, endocrine, and exocrine compartments. These patterns are consistent with the expected organization of pancreatic tumor tissue and indicate that sigmma learns fine-grained microenvironmental structure directly from spatial molecular context. Finally, we asked whether the learned embeddings capture relevant molecular signatures. Differential expression analysis between tumor-associated microenvironments and the immune microenvironment excluded from the tumor identified two chemokines, CXCL2 and CXCL6, both implicated in the recruitment of anti-inflammatory and immunosuppressive myeloid cells that can mediate T cell and B cell exclusion in other cancer contexts [hu2021regulation]. Spatial projection of CXCL2 and CXCL6 expression confirmed that both signals localize sharply to tumor regions in the PAAD section, consistent with a tumor-specific chemokine program (Fig. 4c), Overall, these results show that sigmma not only recovers structural hallmarks of immune exclusion in PAAD but also resolves molecular features that may contribute to the establishment of immunosuppressive cell states within the tumor microenvironment. Thus, sigmma captures biologically interpretable tissue organization across scales and reveals spatially coherent molecular programs that align with the underlying architecture of pancreatic cancer.

### v2

@@ 6 Experiments
We conduct extensive benchmarking on the HEST-1k dataset [jaume2024hest], the largest publicly available dataset of paired HE and ST data. Four subsets of the dataset provide paired HE–Xenium ST data, covering four cancer types: Invasive Ductal Carcinoma (IDC), Pancreatic Adenocarcinoma (PAAD), Skin Cutaneous Melanoma (SKCM), and Lung Adenocarcinoma (LUAD). In addition to the public datasets, we include an in-house skin dataset. Following the tiling scheme in multiple histopathology image foundation models [chen2024towards, vorontsov2024foundation, hoptimus0], we tessellate each WSI into $256\times 256$ pixel-sized tiles at 20x magnification level, which corresponds to $0.5\mu\text{m/pixel}$ resolution. For more details on data preprocessing/data splits, see Suppl. Sec. A.
@@
We comprehensively compare sigmma against three categories of baselines: (1) Uni-modal HE image encoder, UNI [chen2024towards]; (2) Multi-modal vision–language (VL) models pre-trained on natural image–caption pairs, CLIP [radford2021learning], or medical text, PLIP [huang2023visual]; (3) HE–ST contrastive alignment models, including OmiCLIP [chen2025visual], which uses a text encoder by representing ST as a 1D sequence of gene names, and BLEEP [xie2023spatially], a simple MLP operating on batch corrected ST principal components. To ensure fair comparison, all baselines are fine-tuned on the datasets used in this study. For each method, we adopt the hyperparameters reported in the original paper; when unavailable, we determine them through grid search. For more details on the experiment setting, see Suppl. Sec. B.
@@
Here, our focus is to evaluate the quality of the learned image representation for gene expression prediction. Therefore, to avoid introducing biases from different methods’ gene expression decoders, we use the image embedding output by each method and apply a ridge regression for gene expression prediction for each method. Multi-modal alignment consistently enriches HE image embeddings by incorporating ST information (Suppl. Tab. 6). When applying sigmma on top of ResNet50 [he2016deep], H-Optimus-0 [hoptimus0], or UNI [chen2024towards], Sigmma improves representations across backbones, achieving up to 67% lower MSE and 56% higher PCC with UNI. Given its strong gains, we used UNI as the HE encoder backbone for all subsequent experiments.
@@ 6.4 Ablation study
As shown in Tab. 3, we analyze the impact of the core components of sigmma: (1) cell graph, (2) multi-scale loss, and (3) graph sparsification via stochastic edge addition. For the ablation of the cell graph, we replaced the spatial graph representation with a 1D sequence of genes. For the ablation of multi-scale loss, we removed the micro- and meso-scale objectives and trained only with the macro-scale (single-scale) alignment loss. For the ablation of graph sparsification module, instead of selectively sampling edges through stochastic addition, we connected all nodes within neighboring patches, resulting in a fully connected local graph. Tab. 3 shows that each component contributes to performance gains: adding the multi-scale loss and graph sparsification progressively improves prediction accuracy, with all components combined achieving the lowest MSE and highest PCC. The ablation study for Task 2 is provided in the Suppl. Sec. C. Overall, the ablation study shows that each component contributes to performance, with multi-scale loss and graph sparsification having the largest impact.
@@ 6.5 Biological application
We next evaluated whether the multiscale cell-level embeddings learned by sigmma capture meaningful biological structure in a publicly available section from PAAD tissue. For more details on the section, see Suppl. Sec. A, which includes a full description of the dataset composition and sources. We performed clustering of sigmma embeddings at the micro, meso, and macro-scale and projected these clusters back onto a PAAD tissue section, comparing them with reference cell labels (Fig. 4a). Across all three scales, the inferred microenvironments formed coherent and spatially contiguous domains. Most notably at the micro- and meso-scales, sigmma cleanly delineated tumor nests from surrounding non-tumor tissue, despite the absence of any supervision from cell-type labels. To characterize the cellular context represented by these embeddings, we quantified the cell-type composition of each microenvironment cluster, focusing on the micro-scale (Fig. 4b). We identified six resolvable microenvironments. Two microenvironments were composed predominantly of tumor cells. A third represented an inflammatory infiltrate enriched for multiple immune cell types, including T cells and B cells, that were spatially excluded from the tumor region. Three additional microenvironments corresponding to known pancreatic structures, including perivascular, endocrine, and exocrine compartments. These patterns are consistent with the expected organization of pancreatic tumor tissue and indicate that sigmma learns fine-grained microenvironmental structure directly from spatial molecular context. Finally, we asked whether the learned embeddings capture relevant molecular signatures. Differential expression analysis between tumor-associated microenvironments and the immune microenvironment excluded from the tumor identified two chemokines, CXCL2 and CXCL6, both implicated in the recruitment of anti-inflammatory and immunosuppressive myeloid cells that can mediate T cell and B cell exclusion in other cancer contexts [hu2021regulation]. Spatial projection of CXCL2 and CXCL6 expression confirmed that both signals localize sharply to tumor regions in the PAAD section, consistent with a tumor-specific chemokine program (Fig. 4c), Overall, these results show that sigmma not only recovers structural hallmarks of immune exclusion in PAAD but also resolves molecular features that may contribute to the establishment of immunosuppressive cell states within the tumor microenvironment. Thus, sigmma captures biologically interpretable tissue organization across scales and reveals spatially coherent molecular programs that align with the underlying architecture of pancreatic cancer.
@@ 7 Conclusion
8 Author contribution
D.Jeong led the project, proposed the methodology, and designed and executed the core experiments. AV co-designed the experiments, provided project feedback, and supported project management. CRS handled data preprocessing and D.Jafree biological interpretation. MM, KL, and MV assisted with baseline benchmark experiments. SB co-designed and executed a key evaluation task. YC and AB developed and ran the data preprocessing pipeline. LS, VBMS, and ARF generated the datasets. MH and ML supervised the project.
[…]
Supplementary Material
Details on the dataset are provided in Sec. A, and methodological descriptions are included in Sec. B. Additional results that could not be presented in the main text due to space constraints are shown in Sec. C. Finally, Sec. D presents an analysis of challenging cases, limitations, and directions for future work.
Appendix A Data
A.1 Data acquisition
HEST1k dataset.
The HEST1k dataset [jaume2024transcriptomics] is a publicly available benchmark comprising paired HE image and Xenium-ST data, accessible at https://huggingface.co/datasets/MahmoodLab/hest. Table 4 provides a comprehensive summary of all Xenium sections within HEST1k, including tissue types, cell counts, and the presence of cell-type annotations. All sections in the table have paired post-Xenium HE morphology images. Cell-type annotations were derived from the TENX116 section, which serves as the reference dataset for the biological analyses presented in Sec . 6.5.
Table 4: Summary of HEST1k HE-Xenium ST sections.
Tissue type Section ID # of cells Cell-type annotations Source
LUAD TENX118 162,254 Public
TENX141 161,000 Public
PAAD TENX116 190,965 O Public
TENX140 235,099 Public
TENX126 140,702 Public
SKCM TENX115 106,980 Public
TENX117 87,499 Public
IDC TENX95 574,852 Public
TENX99 892,966 Public
NCBI783 142,272 Public
NCBI785 167,780 Public
In-house dataset.
[…]
A.2 Data preprocessing
HE image.
Since the HE image and Xenium ST profile are acquired on different imaging systems, their spatial coordinates are not directly comparable. To place morphological information from the HE image into the same spatial coordinate system, whole-slide HE image and immunofluorescence (IF) were registered using Palom [Chen_PALOM_-_Piecewise_2025], a piecewise registration framework for layers of mosaics. Image registration was performed using the DAPI (IF) and green (HE) channels. Coarse affine alignment was initialized using 4,000 keypoints, followed by local shift refinement and constraint optimization. The final HE image was reconstructed using a blockwise affine transformation and rescaled to $0.5\mu\text{m}$/pixel at 20x magnification level. The Public HEST1k dataset was provided with registration already completed. All sections are tessellated into $256\times 256$ pixel-sized tiles.
ST Xenium profile.
All Xenium ST data is processed with the 10X platform [janesick2023high], which provides built-in cell segmentation and outputs cell-resolved features by default. We extracted ST tiles corresponding to the same spatial region as each HE tile and constructed tile-level representations from the transcriptome profiles of all cells within each tile. For graph representation of each ST tile, we constructed cell graphs using squidpy.gr.spatial_neighbors with default parameters [squidpy_spatial_neighbors], which builds a 6-nearest-neighbor graph from Euclidean coordinates to capture local spatial proximity among cells. For sequence representation of ST in the VL model, we followed Loki preprocessing pipeline [chen2025visual], selecting the top 50 genes by tile-level mean expression and ordering the gene names in descending mean expression.
A.3 Data splitting strategy
HEST1k Dataset.
Due to the limited number of paired HE–Xenium ST sections in public datasets (Tab. 4), section-level split is not feasible, as it would not provide enough data for stable model training. Additionally, assigning spatially adjacent tiles to different splits can lead to spatial leakage across data subsets, since such tiles often share morphological and molecular characteristics. To address these issues, we employ a spatially-stratified tile split for the HEST1k dataset, assigning validation and test tiles to non-overlapping spatial regions (Fig. 5). All splits use an 8:1:1 ratio for training, validation, and testing.
Figure 5: Train-validation-test split of HEST1k dataset.
In-house dataset.
Given the larger number of available sections, we employ a section-level split for the in-house dataset. Each section is assigned entirely to one subset to avoid information leakage across data splits. The dataset is randomly partitioned into training, validation, and test sets with an 8:1:1 ratio.
Appendix B Method
B.1 Implementation detail
We use UNI [chen2024towards] as the HE encoder backbone, which is a ViT-L/16 architecture. For each HE tile, we generate multi-scale crops and resize each patch to the ViT input resolution, 224 $\times$ 224 pixels. Each patch is encoded using the UNI model, and the CLS token output is used as its patch-level embedding. Patch embeddings within each scale are then mean-pooled and L2-normalized to produce a single representation per scale. This procedure is applied from the micro to macro scale, yielding a set of multi-scale HE embeddings.
For the ST encoder. node embeddings are initialized with STEMO features [seb2025STEMO], followed by SAGEConv[hamilton2017inductive]-based message passing and stochastic edge addition layer [piao2022sparse]. All GNN components are implemented using the dgl library. Graph-level embeddings are then obtained via attention pooling. This procedure is applied hierarchically from micro to macro scales to produce multi-scale ST representations.
At each scale, HE and ST embeddings are passed through simple MLP projection heads to obtain vectors of the same dimensionality, and a symmetric InfoNCE loss is then applied to align the modalities. The hidden dimensions for HE and ST embeddings are set to 256 or 512. All projection and encoder layers use LeakyReLU activations.
We optimize the model using AdamW with a StepLR scheduler and weight decay. To obtain a larger effective batch size, which is essential for contrastive learning, we employ 10 steps of gradient accumulation with a batch size of 128 and incorporate cross-batch negatives. All experiments are conducted on a single A100 SXM4 GPU with 80GB RAM.
[…]
B.2 Evaluation metrics.
Gene expression prediction.
For each tile, we assess how accurately the model predicts the expression vector over the top $G$ (G=50) highly variable genes. Let $y_{i,g}$ and $\hat{y}_{i,g}$ denote the ground-truth and predicted expression values of gene $g$ in tile $i$. Tile-level MSE and PCC are computed across genes as follows:
$\text{MSE}_{i}$ $=\frac{1}{G}\sum_{g=1}^{G}(y_{i,g}-\hat{y}_{i,g})^{2}$ (1)
$\text{PCC}_{i}$ $=\frac{(\mathbf{y}_{i}-\bar{\mathbf{y}}_{i})^{\top}(\hat{\mathbf{y}}_{i}-\bar{\hat{\mathbf{y}}}_{i})}{\|\mathbf{y}_{i}-\bar{\mathbf{y}}_{i}\|_{2}\;\|\hat{\mathbf{y}}_{i}-\bar{\hat{\mathbf{y}}}_{i}\|_{2}}$ (2)
We report dataset-level performance as the mean $\pm$ standard deviation of $\text{MSE}_{i}$ and $\text{PCC}_{i}$ across all tiles.
Cross-modal retrieval.
For a query tile $q$ from one modality (HE or ST), we rank all tiles in the other modality by embedding similarity. Retrieval accuracy is measured by whether the paired tile appears within the top-pp%. Recall@pp% is defined as follows:
$\text{Recall@}p\%=\frac{1}{N}\sum_{q=1}^{N}\mathbbm{1}\!\left(y_{q}\in\text{TopK}(q)\right),$ (3)
where $N$ denotes the number of query tiles, $K=\lfloor p\%\times N\rfloor$ and $\mathbbm{1}(\cdot)$ is the indicator function.
B.3 Fine-tuning of baselines.
All baselines were either fine-tuned on our standardized data splits or evaluated using their official checkpoints. Each model was fine-tuned using the loss function specified in the original implementation. For BLEEP, no pretrained BLEEP checkpoint is provided, and only the ResNet50 [he2016deep] backbone is publicly available; therefore, we initialized BLEEP with this backbone and fine-tuned the model on our dataset to ensure a consistent and fair comparison.
Appendix C Result
C.1 Ablation study
Here, we present ablation results for the core components of the cross-retrieval task (Task 2) on the HEST1k-LUAD dataset. As shown in Tab. 5, each component contributes to performance gains in the retrieval task: adding the multi-scale loss and graph sparsification progressively improves retrieval accuracy, and integrating all components yields the best overall performance.
Table 5: Ablation of core components of sigmma on HEST1k-LUAD dataset for the cross-modal retrieval task.
Components Task 2. HE $\rightarrow$ ST
Cell graph Multi-scale loss Graph sparsification R@5% R@10% R@15%
0.517 0.694 0.768
0.480 0.639 0.737
0.550 0.667 0.786
0.590 0.728 0.826
Components Task 2. ST $\rightarrow$ HE
Cell graph Multi-scale loss Graph sparsification R@5% R@10% R@15%
0.529 0.673 0.780
0.459 0.621 0.722
0.514 0.685 0.761
0.602 0.768 0.813
Extending the ablation in the main manuscript, we additionally explored (1) different HE image backbones, ResNet50 [he2016deep], H-Optimus-0 [hoptimus0], and UNI [chen2024towards], and (2) different ST backbones for cell embedding initialization, Harmony [korsunsky2019fast], Novae [blampeyNovae2024], and STEMO [seb2025STEMO]. These experiments assess the impact of backbone choices and support the architectural decisions made in our final model. Sigmma, built upon the UNI vision backbone, consistently achieves the best performance across both downstream tasks (Tab. 6, Tab. 7). On the other hand, among the ST backbones, STEMO tended to perform well overall across both downstream tasks (Tab. 8, Tab. 9). Accordingly, these results justify selecting UNI and STEMO as the vision and ST backbones in our framework.
Table 6: Ablation of vision backbone in HEST1k-LUAD dataset for task 1.
Task 1. Gene expression prediction.
Model MSE (↓) PCC (↑)
[…]
Table 7: Ablation of vision backbones in HEST1k-LUAD dataset for task 2.
Task 2. Cross-modal retrieval.
HE $\rightarrow$ ST ST $\rightarrow$ HE
Model R@5% R@10% R@15% R@5% R@10% R@15%
Sigmma (ResNet50) 0.086 0.199 0.260 0.135 0.229 0.333
Sigmma (H-Optimus-0) 0.563 0.676 0.777 0.554 0.688 0.774
Sigmma (UNI) 0.590 0.728 0.826 0.602 0.768 0.813
Table 8: Ablation of ST backbone in HEST1k-LUAD dataset for task 1.
Task 1. Gene expression prediction.
Model MSE (↓) PCC (↑)
[…]
Sigmma (Novae) $1±\pm0.0$03 $6±\pm0.0$48
[…]
Table 9: Ablation of ST backbone in HEST1k-LUAD dataset for task 2.
Task 2. Cross-modal retrieval.
HE $\rightarrow$ ST ST $\rightarrow$ HE
Model R@5% R@10% R@15% R@5% R@10% R@15%
Sigmma (Harmony) 0.529 0.723 0.820 0.526 0.730 0.830
Sigmma (Novae) 0.039 0.086 0.148 0.049 0.099 0.148
Sigmma (STEMO) 0.590 0.728 0.826 0.602 0.768 0.813
C.2 Sensitivity analysis on tile size
We further conduct a sensitivity analysis on tile size, as tile resolution is a common source of variability in histopathology pipelines. Tile size 224 is a common choice in general vision models, while 256 is the standard image size used in many histopathology foundation models [chen2024towards, vorontsov2024foundation, hoptimus0]. As shown in Tab. 10 and Tab. 11, performance varies moderately across 224, 256, and 512 px tiles, but sigmma shows no collapse or strong dependence on any specific configuration. Mid-sized tiles (224/256 px) perform well across both tasks, while 512 px tiles show reduced performance, potentially due to increased heterogeneity within larger regions. Overall, this sensitivity analysis shows that while performance varies with tile resolution, sigmma remains generally robust, and medium tile sizes tend to provide favorable performance across tasks.
Table 10: Sensitivity analysis for task 1.
Task 1. Gene expression prediction.
Model MSE (↓) PCC (↑)
[…]
Table 11: Sensitivity analysis for task 2.
Task 2. Cross-modal retrieval.
HE $\rightarrow$ ST ST $\rightarrow$ HE
Model R@5% R@10% R@15% R@5% R@10% R@15%
224 $\times$ 224 0.616 0.755 0.817 0.598 0.744 0.794
256 $\times$ 256 0.590 0.728 0.826 0.602 0.768 0.813
512 $\times$ 512 0.453 0.605 0.721 0.523 0.640 0.733
Appendix D Discussions
D.1 Analysis of challenging cases
Understanding the PCC–MSE discrepancy.
Although sigmma achieves high PCC across genes, indicating that it accurately captures the relative variation of expression across tiles, inspection of calibration plots reveals a consistent miscalibration in absolute prediction values. As shown in Fig. 6, the regression line has slopes << 1 and positive intercepts, indicating that the model underestimates variation while introducing a systematic bias. This global calibration mismatch increases MSE despite preserving rank-order consistency, explaining the discrepancy between PCC and MSE observed in Tab. 1.
Figure 6: Calibration analysis of gene expression predictions.
Embedding similarity challenges IDC retrieval.
[…]
Figure 7: Cosine similarity of HE embeddings.
D.2 Limitations
Although sigmma learns meaningful multi-modal representation, several limitations remain. The model’s generalizability is constrained by the limited range and diversity of available paired HE–Xenium ST datasets, hindering robust performance across heterogeneous tissue types. Additionally, because the approach relies on hierarchical spatial graphs constructed from cell segmentation, its effectiveness is inherently dependent on segmentation quality, which may be variable in complex tissue contexts.
D.3 Future work
Future extensions of sigmma include evaluating the framework on other single-cell–resolution spatial transcriptomics platforms (e.g., CosMx, MERSCOPE) to assess cross-platform robustness. In addition, sigmma’s hierarchical design naturally scales to the WSI-level task, which we plan to explore as a next step. Finally, extending retrieval evaluation beyond within-tissue settings to cross-tissue scenarios may reveal how well the learned representations generalize across distinct morphological and molecular contexts.